# intersection.bib

### v2

@@ 4 Bounds for number of cars
In this section we prove the bounds for the tail distribution of $N$. In particular, we will prove there exist $c_{1},C_{1}>0$ (depending on $\varepsilon$) such that for all $n$ sufficiently large
@@
$N_{v}:=|\{u\in\mathbb{Z}^{2}:v\in\Gamma_{u}^{\theta_{u}}\}|$
@@
$\widetilde{N}:=|\{u\in\mathbb{Z}^{2}:-2n\leq\phi(u)\leq 0\text{ and }\exists v\in V\text{ such that }v\in\Gamma_{u}^{\theta_{u}}\}|.$
@@
Same as before, for $1\leq i\leq\lfloor n^{1/3}\rfloor$ and for a uniformly bounded sequence $\{c_{n}\}$ we partition the interval $(\epsilon,\frac{\pi}{2}-\epsilon$) into intervals $A_{i}$ each of equal length $\frac{c_{n}(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$. For $1\leq i\leq\lfloor n^{1/3}\rfloor$ we define the following random variables.
$\widetilde{N}_{i}:=|\{u\in\mathbb{Z}^{2}:-2n\leq\phi(u)\leq 0,\theta_{u}\in A_{i}\text{ and }\exists v\in V\text{ such that }v\in\Gamma_{u}^{\theta_{u}}\}|$
@@
$\{\widetilde{N}\geq\ell n^{5/3}\log n\}\subset\bigcup_{i=1}^{\lfloor n^{1/3}\rfloor}\{\widetilde{N}_{i}\geq\ell n^{4/3}\log n\}.$
We will show that for each $1\leq i\leq\lfloor n^{1/3}\rfloor$ there exist $C,c>0$ (depending only on $\epsilon$) such that for sufficiently large $n$ (depending only on $\epsilon$) and $\ell<n^{1/6}$
$\mathbb{P}(\widetilde{N}_{i}\geq\ell n^{4/3})\leq Ce^{-c\ell}.$ (4.1)
@@
Figure 6: To prove the upper bound in Theorem 2.4 we fix $1\leq i\leq\lfloor n^{1/3}\rfloor$. The event that there are more than $\ell n^{4/3}$ many $u$ with $\theta_{u}\in A_{i}$ and $\Gamma_{u}^{\theta_{u}}$ intersecting $V$ can happen in two ways. Either the above $u$ lies outside $\mathsf{R}_{i}$. For sufficiently large $\ell$ this will imply large transversal fluctuation for either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$. Proposition 3.3 gives a stretched exponential upper bound for this event. The other possibility is there are more than $\ell n^{4/3}$ many vertices inside $\mathsf{R}_{i}$ that choose an angle in $A_{i}$. Using Hoeffding’s inequality we get a stretched exponential upper bound for this event. Combining the above we get (4.1).
@@
  • $\mathcal{H}_{i}:=\{\exists u\in\mathbb{Z}^{2}\setminus\mathsf{R}_{i}$ such that $\theta_{u}\in A_{i}$ and $v\in\Gamma_{u}$ for some $v\in V$}.
@@
$\mathbb{P}(\widetilde{N}_{i}\geq\ell n^{4/3})\leq\mathbb{P}(\mathcal{H}_{i})+\mathbb{P}(\{\widetilde{N}_{i}\geq\ell n^{4/3}\}\cap(\mathcal{H}_{i})^{c}).$ (4.2)
Note that by planarity the event $\mathcal{H}_{i}$ can happen in two ways. Either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will leave $\mathsf{R}_{i}$ or they will intersect $V$ (see Figure 6). Since, $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}}$, for sufficiently large $\ell$ (depending on $\epsilon$), the event $\mathcal{H}_{i}$ will imply that either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will have transversal fluctuation larger than $\frac{\ell}{64}n^{2/3}$. We have the following proposition.
Hence, for sufficiently large $\ell$
@@
for some $C,c>0$ depending on $\epsilon.$ We consider the second term in (4.2). On the event $\{\widetilde{N}_{i}\geq\ell n^{4/3}\}\cap(\mathcal{H}_{i})^{c}$ there are at least $\ell n^{4/3}$ many vertices inside $\mathsf{R}_{i}$ that chooses an angle from $A_{i}$. For large $\ell$ and $n$ there are at most $\frac{\ell}{4}n^{5/3}$ many vertices inside $\mathsf{R}_{i}$. Consider the collection of i.i.d. random variables $\{\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\}_{u\in\mathsf{R}_{i}}$ with mean $\frac{c_{n}}{n^{1/3}}$. So,
$\{\widetilde{N}_{i}\geq\ell n^{4/3}\}\cap(\mathcal{H}_{i})^{c}\subset\{\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\geq\ell n^{4/3}\}.$
@@
($\mathbb{P}\left(\{\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\geq\ell n^{4/3}\}\right)\leq\mathbb{P}\left(\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}-\mathbb{E}(\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}})\geq\frac{\ell}{2}n^{4/3}\right)\leq e^{-c\ell}.$
@@
$\mathbb{P}(\widetilde{N}_{i}\geq\ell n^{4/3}\log n)\leq Ce^{-c\ell\log n}.$
@@
$\mathbb{P}(\widehat{N}\geq\ell n^{1/3}\log n)\leq C^{\prime}e^{-c^{\prime}\ell}.$
So, we have there exists $\ell_{0}$ (depending only on $\epsilon$) and $\widetilde{C}$ (depending only on $\epsilon$)
@@
Same as before, first we fix large constants $M$ and $\ell_{0}$ which will be chosen later. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$, we consider the intervals $A_{i}\subset(\epsilon,\frac{\pi}{2}-\epsilon)$, each of length $\frac{(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$ and midpoint $\frac{Mi(\frac{\pi}{2}-2\epsilon)}{n^{1/3}}$. Corresponding to each $A_{i}$ we define the following random variables.
$N_{i}:=|\{u\in\mathbb{Z}^{2}:\theta_{u}\in A_{i}\text{ and }\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\}|.$
@@
$\bigcup_{i=1}^{[\frac{n^{1/3}}{2M}]}\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\}\subset\{N\geq\frac{n^{4/3}}{\ell_{0}}\}.$
@@
$\mathcal{I}_{i}:=\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\}.$
Applying inclusion exclusion principle on the union of $\mathcal{I}_{i}$ and considering the first two terms we have the following lower bound.
$\mathbb{P}(N\geq\frac{n^{4/3}}{\ell_{0}})\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{I}_{i})-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{I}_{i}\cap\mathcal{I}_{j}).$ (4.3)
@@
$N_{i}^{v}:=|\{u\in\mathbb{Z}^{2}:\theta_{u}\in A_{i}\text{ and }v\in\Gamma_{u}^{\theta_{u}}\}.$
$\lceil 4Mn^{2/3}\rceil\mathbb{P}\left(\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\}\right)=\sum_{v\in V_{M}}\mathbb{P}\left(\{N_{i}^{v}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D_{v}\leq 2n\}\right)=\mathbb{E}(\widehat{N_{i}}),$
@@
$\widehat{N_{i}}:=\sum_{v\in V_{M}}\mathbbm{1}_{\{N_{i}^{v}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D_{v}\leq 2n\}}.$
Figure 7: To find lower bound for the first sum in (4.3) we fix an $i$. Using Proposition 3.3 we can fix $M$ large enough so that with high probability $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ are contained in $\mathsf{R}_{i}^{M}$ and do not intersect $\mathsf{R}_{i}$. On this event any car starting from $\mathsf{R}_{i}$ and taking angle in $A_{i}$ will intersect $\mathcal{L}_{0}$ on $V_{M}$. Further, using [BHS18, Theorem 3.10] depending on this $M$ we can chose $\ell_{0}$ large enough so that with large probability there are at most $\ell_{0}$ many distinct points on $V_{M}$ carrying a geodesic starting from $V_{M}^{\prime}$ and ending at $V_{M}^{\prime\prime}$. Also, with positive probability there are at least $n^{4/3}$ any vertices $u\in\mathsf{R}_{i}$ such that $\theta_{u}\in A_{i}$. The last event is independent of the first two events. Hence, combining all these we get on a positive probability event there exists $v\in V_{M}$ such that there are at least $\frac{n^{4/3}}{\ell_{0}}$ many $u$ with $\theta_{u}\in A_{i}$ and $v\in\Gamma_{u}^{\theta_{u}}$. Now, applying an averaging argument and summing over all $i$ give a lower bound for the first sum in (4.3).
@@
  • $\mathsf{TF}_{i}$ is the event that $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ are contained within $\mathsf{R}_{i}^{M}$ and $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ do not enter $\mathsf{R}_{i}$.
As $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}}$, the event $\mathsf{TF}_{i}$ will imply for sufficiently large $M$ either $\Gamma_{v_{1}}^{\beta_{i}}$ or $\Gamma_{v_{2}}^{\alpha_{i}}$ will have transversal fluctuation more than $\frac{M}{4}n^{2/3}$. Using Proposition 3.3 we can chose $M$ (depending only on $\epsilon$) large enough and $n$ sufficiently large enough (depending on $M$) so that the following happens.
@@
On $\mathsf{TF}_{i}$, for any $u\in\mathsf{R}_{i}$ with $\theta_{u}\in A_{i},v\in\Gamma_{u}^{\theta_{u}}$, for some $v\in V_{M}.$ Now we define an equivalence class. Let $n_{i}^{\prime\prime}$ denote the intersection point of $y=\tan(\theta_{i})x$ with $\mathcal{L}_{n}$. Consider the two line segments $V_{M}^{\prime}$ (resp. $V_{M}^{\prime\prime}$) with $\lceil 4Mn^{2/3}\rceil$ vertices on $\mathcal{L}_{-n}$ (resp. $\mathcal{L}_{n}$) with midpoint $n_{i}^{\prime}$ (resp. $n_{i}^{\prime\prime}$). For $u,v\in V_{M}^{\prime}$ and $u^{\prime},v^{\prime}\in V_{M}^{\prime\prime},$ we say $(u,v)\sim(u^{\prime},v^{\prime})$ if $\Gamma_{u,v}(0)=\Gamma_{u^{\prime},v^{\prime}}(0).$ Let $E_{i}$ denote the number of equivalence classes. Using [BHS18, Theorem 3.10], we can chose $\ell_{0}$ large enough (depending on $M$) so that for all sufficiently large $n$
$\mathbb{P}(E_{i}\leq\ell_{0})\geq 0.99.$
@@
Note that for large $n$, $\mathbb{E}(Y_{n})\geq\frac{n^{4/3}}{4}$ and $\operatorname{{\mathbb{V}ar}}(Y_{n})=c_{n}(n^{4/3}-n)$ for some uniformly bounded sequence $\{c_{n}\}.$ By Berry-Esseen inequality we have the following for all $n$ and a constant $c$
@@
($\mathbb{P}(\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\})\geq\frac{1}{4}.$
@@
$\mathsf{TF}_{i}\cap\{E_{i}\leq\ell_{0}\}\cap\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\}\subset\{\widehat{N_{i}}\geq 1\}.$ (4.4)
Further, the event $\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\}$ is independent of $\mathsf{TF}_{i}\cap\{E_{i}\leq\ell_{0}\}$. So, combining all the above for sufficiently large $n,$ there exists $c_{1}>0$ ($c_{1}$ is a fixed constant, does not depend on anything) such that
$\mathbb{E}(\widehat{N_{i}})\geq\mathbb{P}\left(\mathsf{TF}_{i}\cap\{E_{i}\leq\ell_{0}\}\cap\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\}\right)\geq c_{1}.$
@@
($\mathbb{P}\left(\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\}\right)\geq\frac{c_{1}}{4Mn^{2/3}}.$ (4.5)

### v3

@@ 4 Bounds for number of cars
In this section we prove the bounds for the tail distribution of $N$. Recall that $N$ was defined to be the number of cars that pass through the origin, and $D$ was the greatest distance a car travelled to the origin. More precisely,
$N:$ =#$=\#\left\{u\in\mathbb{Z}^{2}:\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\right\},$
In particular, we will prove there exist $c_{1},C_{1}>0$ (depending on $\varepsilon$) such that for all $n$ sufficiently large
@@
$N_{v}:=\#\left\{u\in\mathbb{Z}^{2}:v\in\Gamma_{u}^{\theta_{u}}\right\}.$
@@
$\widetilde{N}:=\#\{u\in\mathbb{Z}^{2}:-2n\leq\phi(u)\leq 0\text{ and }\exists v\in V\text{ such that }v\in\Gamma_{u}^{\theta_{u}}\}.$
@@
Same as before, for $1\leq i\leq\lfloor n^{1/3}\rfloor$ and for $c_{n}=\frac{n^{1/3}}{\lfloor n^{1/3}\rfloor}$ we partition the interval $(\varepsilon,\frac{\pi}{2}-\varepsilon$) into intervals $A_{i}$ each of equal length $\frac{c_{n}(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$. For $1\leq i\leq\lfloor n^{1/3}\rfloor$ we define the following random variables.
$\widetilde{N}_{i}:=\#\left\{u\in\mathbb{Z}^{2}:-2n\leq\phi(u)\leq 0,\theta_{u}\in A_{i}\text{ and }\exists v\in V\text{ such that }v\in\Gamma_{u}^{\theta_{u}}\right\}.$
@@
$\left\{\widetilde{N}\geq\ell n^{5/3}\log n\right\}\subset\bigcup_{i=1}^{\lfloor n^{1/3}\rfloor}\left\{\widetilde{N}_{i}\geq\ell n^{4/3}\log n\right\}.$
We will show that for each $1\leq i\leq\lfloor n^{1/3}\rfloor$ there exist $C,c>0$ (depending only on $\varepsilon$) such that for sufficiently large $n$ (depending only on $\varepsilon$) and $\ell<n^{1/6}$
$\mathbb{P}\left(\widetilde{N}_{i}\geq\ell n^{4/3}\right)\leq Ce^{-c\ell}.$ (4.1)
@@
Figure 7: To prove the upper bound in Theorem 2.4 we fix $1\leq i\leq\lfloor n^{1/3}\rfloor$. The event that there are more than $\ell n^{4/3}$ many $u$ with $\theta_{u}\in A_{i}$ and $\Gamma_{u}^{\theta_{u}}$ intersecting $V$ can happen in two ways. Either the above $u$ lies outside $\mathsf{R}_{i}$. For sufficiently large $\ell$ this will imply large transversal fluctuation for either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ (the case when $\Gamma^{\alpha_{i}}_{v_{1}}$ will have large transversal fluctuation is shown in the figure). Proposition 6.3 gives a stretched exponential upper bound for this event. The other possibility is that there are more than $\ell n^{4/3}$ many vertices inside $\mathsf{R}_{i}$ that choose an angle in $A_{i}$. Using Hoeffding’s inequality we get a stretched exponential upper bound for this event. Combining the above we get (4.1).
@@
  • $\mathcal{H}_{i}:=\left\{\exists u\in\mathbb{Z}^{2}\setminus\mathsf{R}_{i}\text{ such that }\theta_{u}\in A_{i}\text{ and }v\in\Gamma^{\theta_{u}}_{u}\text{ for some }v\in V\right\}$.
@@
$\mathbb{P}\left(\widetilde{N}_{i}\geq\ell n^{4/3}\right)\leq\mathbb{P}(\mathcal{H}_{i})+\mathbb{P}\left(\left\{\widetilde{N}_{i}\geq\ell n^{4/3}\right\}\cap(\mathcal{H}_{i})^{c}\right).$ (4.2)
Note that by planarity the event $\mathcal{H}_{i}$ can happen in two ways. Either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will leave $\mathsf{R}_{i}$ or they will have large transversal fluctuation from the straight lines starting from $v_{1}$ in the direction $\alpha_{i}$ and starting from $v_{2}$ in the direction $\beta_{i}$ respectively. (see Figure 7). Since, $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}}$, for sufficiently large $\ell$ (depending on $\varepsilon$), the event $\mathcal{H}_{i}$ will imply that either $\Gamma_{v_{1}}^{\alpha_{i}}$ or $\Gamma_{v_{2}}^{\beta_{i}}$ will have transversal fluctuation larger than $\frac{\ell}{64}n^{2/3}$. By Proposition 6.3 below we have Hence, for sufficiently large $\ell$
@@
for some $C,c>0$ depending on $\varepsilon.$ We consider the second term in (4.2). On the event $\left\{\widetilde{N}_{i}\geq\ell n^{4/3}\right\}\cap(\mathcal{H}_{i})^{c}$ there are at least $\ell n^{4/3}$ many vertices inside $\mathsf{R}_{i}$ that chooses an angle from $A_{i}$. For large $\ell$ and $n$ there are at most $\frac{\ell}{4}n^{5/3}$ many vertices inside $\mathsf{R}_{i}$. Consider the collection of i.i.d. random variables $\{\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\}_{u\in\mathsf{R}_{i}}$ with mean $\frac{c_{n}}{n^{1/3}}$. So,
$\{\widetilde{N}_{i}\geq\ell n^{4/3}\}\cap(\mathcal{H}_{i})^{c}\subset\left\{\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\geq\ell n^{4/3}\right\}.$
@@
$\mathbb{P}\left(\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\geq\ell n^{4/3}\right)\leq\mathbb{P}\left(\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}-\mathbb{E}\left(\sum_{u\in\mathsf{R}_{i}}\mathbbm{1}_{\{\theta_{u}\in A_{i}\}}\right)\geq\frac{\ell}{2}n^{4/3}\right)\leq e^{-c\ell}.$
@@
$\mathbb{P}\left(\widetilde{N}_{i}\geq\ell n^{4/3}\log n\right)\leq Ce^{-c\ell\log n}.$
@@
$\mathbb{P}\left(\widehat{N}\geq\ell n^{1/3}\log n\right)\leq C^{\prime}e^{-c^{\prime}\ell}.$
So, we have that there exists $\ell_{0}$ (depending only on $\varepsilon$) and $\widetilde{C}$ (depending only on $\varepsilon$)
@@
Same as before, first we fix large constants $M$ and $\ell_{0}$ which will be chosen later. For $1\leq i\leq[\frac{n^{1/3}}{2M}]$, we consider the intervals $A_{i}\subset(\varepsilon,\frac{\pi}{2}-\varepsilon)$, each of length $\frac{(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$ and midpoint $\frac{Mi(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$. Corresponding to each $A_{i}$ we define the following random variables.
$N_{i}:=\#\left\{u\in\mathbb{Z}^{2}:\theta_{u}\in A_{i}\text{ and }\boldsymbol{0}\in\Gamma_{u}^{\theta_{u}}\right\}.$
@@
$\bigcup_{i=1}^{[\frac{n^{1/3}}{2M}]}\left\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\right\}\subset\left\{N\geq\frac{n^{4/3}}{\ell_{0}}\right\}.$
@@
$\mathcal{I}_{i}:=\left\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\right\}.$
Applying inclusion exclusion principle on the union of $\mathcal{I}_{i}$ and considering the first two terms (Proposition 6.1) we have the following lower bound.
$\mathbb{P}\left(N\geq\frac{n^{4/3}}{\ell_{0}}\right)\geq\sum_{i=1}^{[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{I}_{i})-\sum_{1\leq i<j\leq[\frac{n^{1/3}}{2M}]}\mathbb{P}(\mathcal{I}_{i}\cap\mathcal{I}_{j}).$ (4.3)
@@
$N_{i}^{v}:=\#\left\{u\in\mathbb{Z}^{2}:\theta_{u}\in A_{i}\text{ and }v\in\Gamma_{u}^{\theta_{u}}\right\}.$
$\lceil 4Mn^{2/3}\rceil\mathbb{P}\left(\left\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\right\}\right)=\sum_{v\in V_{M}}\mathbb{P}\left(\left\{N_{i}^{v}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D_{v}\leq 2n\right\}\right)$
$=\mathbb{E}(\widehat{N_{i}}),$
@@
$\widehat{N_{i}}:=\sum_{v\in V_{M}}\mathbbm{1}_{\left\{N_{i}^{v}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D_{v}\leq 2n\right\}}.$
Figure 8: To find lower bound for the first sum in (4.3) we fix an $i$. Using Proposition 6.3 we can fix $M$ large enough so that with high probability $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ are contained in $\mathsf{R}_{i}^{M}$ and do not intersect $\mathsf{R}_{i}$. On this event any car starting from $\mathsf{R}_{i}$ and taking angle in $A_{i}$ will intersect $\mathcal{L}_{0}$ on $V_{M}$. Further, using [BHS18, Theorem 3.10] depending on this $M$ we can choose $\ell_{0}$ large enough so that with large probability there are at most $\ell_{0}$ many distinct points on $V_{M}$ carrying a geodesic starting from $V_{M}^{\prime}$ and ending at $V_{M}^{\prime\prime}$. Also, with positive probability there are at least $n^{4/3}$ many vertices $u\in\mathsf{R}_{i}$ such that $\theta_{u}\in A_{i}$. The last event is independent of the first two events. Hence, combining all these we get on a positive probability event there exists $v\in V_{M}$ such that there are at least $\frac{n^{4/3}}{\ell_{0}}$ many $u$ with $\theta_{u}\in A_{i}$ and $v\in\Gamma_{u}^{\theta_{u}}$. Now, applying an averaging argument and summing over all $i$ give a lower bound for the first sum in (4.3).
@@
  • $\mathsf{TF}_{i}$ is the event that $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ are contained between the longer sides of $\mathsf{R}_{i}^{M}$ and $\Gamma_{v_{1}}^{\beta_{i}}$ and $\Gamma_{v_{2}}^{\alpha_{i}}$ do not enter $\mathsf{R}_{i}$.
As $(\beta_{i}-\alpha_{i})$ is of order $\frac{1}{n^{1/3}}$, the event $\mathsf{TF}_{i}$ will imply for sufficiently large $M$ either $\Gamma_{v_{1}}^{\beta_{i}}$ or $\Gamma_{v_{2}}^{\alpha_{i}}$ will have transversal fluctuation more than $\frac{M}{4}n^{2/3}$. Using Proposition 6.3 we can choose $M$ (depending only on $\varepsilon$) large enough and $n$ sufficiently large enough (depending on $M$) so that the following happens.
@@
On $\mathsf{TF}_{i}$, for any $u\in\mathsf{R}_{i}$ with $\theta_{u}\in A_{i},v\in\Gamma_{u}^{\theta_{u}}$, for some $v\in V_{M}.$ Now we define an equivalence class. Let $n_{i}^{\prime\prime}$ denote the intersection point of $y=\tan(\theta_{i})x$ with $\mathcal{L}_{n}$. Consider the two line segments $V_{M}^{\prime}$ (resp. $V_{M}^{\prime\prime}$) with $\lceil 4Mn^{2/3}\rceil$ vertices on $\mathcal{L}_{-n}$ (resp. $\mathcal{L}_{n}$) with midpoint $n_{i}^{\prime}$ (resp. $n_{i}^{\prime\prime}$). For $u,v\in V_{M}^{\prime}$ and $u^{\prime},v^{\prime}\in V_{M}^{\prime\prime},$ we say $(u,v)\sim(u^{\prime},v^{\prime})$ if $\Gamma_{u,v}(0)=\Gamma_{u^{\prime},v^{\prime}}(0).$ Let $E_{i}$ denote the number of equivalence classes. Using Proposition 6.4, we can choose $\ell_{0}$ large enough (depending on $M$) so that for all sufficiently large $n$
$\mathbb{P}\left(E_{i}\leq\frac{\ell_{0}}{4}\right)\geq 0.99.$
@@
Note that for large $n$, $\mathbb{E}(Y_{n})\geq\frac{n^{4/3}}{4}$ and $\operatorname{{\mathbb{V}ar}}(Y_{n})=n\lceil\frac{n^{2/3}}{4}\rceil\left(\frac{1}{n^{1/3}}-\frac{1}{n^{2/3}}\right)$. By the Berry-Esseen inequality (Proposition 6.2) we have the following for all $n$ and a constant $c$
@@
($\mathbb{P}\left(\left\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\right\}\right)\geq\frac{1}{4}.$
@@
$\mathsf{TF}_{i}\cap\left\{E_{i}\leq\frac{\ell_{0}}{4}\right\}\cap\left\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\right\}\subset\{\widehat{N_{i}}\geq 1\}.$ (4.4)
Further, the event $\left\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\right\}$ is independent of $\mathsf{TF}_{i}\cap\left\{E_{i}\leq\frac{\ell_{0}}{4}\right\}$. So, combining all the above for sufficiently large $n,$ there exists $c_{1}>0$ ($c_{1}$ is a fixed constant, does not depend on anything) such that
$\mathbb{E}(\widehat{N_{i}})\geq\mathbb{P}\left(\mathsf{TF}_{i}\cap\left\{E_{i}\leq\frac{\ell_{0}}{4}\right\}\cap\left\{Y_{n}-\frac{n^{4/3}}{4}\geq 0\right\}\right)\geq c_{1}.$
@@
($\mathbb{P}\left(\left\{N_{i}\geq\frac{n^{4/3}}{\ell_{0}},n\leq D\leq 2n\right\}\right)\geq\frac{c_{1}}{4Mn^{2/3}}.$ (4.5)